6. A new rocket-launching system is being considered for deployment of small, short-range rockets. The existing system has p = 0.8 as the probability of a successful launch. A sample of 40 experimental launches is made with the new system, and 34 are successful. Using 95% confidence, the confidence interval is 0.85 + 0.111. What is the sample size required @ 90% level of confidence?


1
Expert's answer
2021-11-22T12:13:06-0500
"\hat{p}=\dfrac{34}{40}=0.85"

"z_c=1.6449"

"SE=z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}}"

"n=\dfrac{z_c^2\hat{p}(1-\hat{p})}{SE^2}"

Given "SE=0.111"


"n=\dfrac{(1.6449)^2(0.85)(1-0.85)}{(0.111)^2}=28"